# Towards Constraint-aware Learning for Resource Allocation in NFV-enabled Networks

### v1

@@ 3.3 Constraint-aware Graph Representation
The VNE processing is governed by complex and multifaceted constraints, presenting challenges in representation learning. These include the interaction of cross-graph status and the assessment of bandwidth-constrained path connectivity. To address this issue, we propose a constraint-aware graph representation method with a heterogeneous modeling module for cross-graph status fusion and a contrastive learning-based module to enhance path connectivity awareness. This method efficiently perceives the complex constraints of VNE, providing a higher feasibility guarantee of solutions.
@@
To address this challenge, we propose a novel path-bandwidth contrast method to enhance bandwidth constraint awareness through contrastive learning, whose core idea is creating augmented views with feasibility-consistency augmentations and making node representations in these views close.
Definition 4 (Feasibility-consistency Augmentations).
@@
These augmentations generate multiple views of the original heterogeneous graph, which maintain the same feasibility semantics before and after their application. Following this principle, we develop several augmentation methods by modifying the topology of either VN or PN without impacting solution feasibility, which are described as follows. (a) Physical Link Addition $\phi^{A}$. We add a specific number $\epsilon\cdot|N^{p}|$ of physical links in PN, whose bandwidth resources are equal to the difference between the smallest requirements among all virtual links and 1, i.e., $\min_{l_{v}\in L_{v}}\mathcal{B}(l_{v})-1$. (b) Virtual Link Addition $\phi^{B}$. We add a specific number $\epsilon\cdot|N^{v}|$ of virtual links that require a zero bandwidth resource to enhance the complexity and connectivity of the VN. Here, $\epsilon$ is an augment ratio that determines the proportion of links to be added based on the number of nodes in the network.
After applying these augmentations, we create two new views $G_{I}^{A}=\phi_{A}(G_{I})$ and $G_{I}^{B}=\phi^{B}(G_{I})$, which have same feasibility semantics of VNE instance $I$. Using our heterogeneous graph network, we extract node representations under views $G_{I}^{A}$ and $G_{I}^{B}$, denoted as $Z^{A}$ and $Z^{B}$, respectively. Subsequently, we utilize contrastive learning to enhance the proximity of node representations under the augmented views $G_{I}^{A}$ and $G_{I}^{B}$. This necessitates that the model precisely discerns the noisy implications of those links with less bandwidth that play no impact on solution feasibility. Through this method, we aim to enhance the model’s sensitivity towards link bandwidth, effectively mitigate the influence of irrelevant links in the GNN propagation process, and bolster its overall awareness of bandwidth constraints. In this work, we adopt the Barlow Twins method (Zbontar et al., 2021) for its simplicity and effectiveness, which circumvents negative sample selection and maintains the original network architecture. Given the embeddings under two augmented views, $H^{a}$ and $H^{b}$, we use this contrastive loss $L_{\textit{CL}}$ to reduce redundancy between embedding components by aligning their cross-correlation matrix with the identity matrix. This unsupervised loss can be seamlessly integrated into the training process of RL. See Appendix D.4 for the details of Barlow Twins method.
@@ 3.4 Computational Complexity Analysis
Note that CONAL solely uses its surrogate policy and path-bandwidth contrast module for policy optimization during training. During inference, CONAL has a computational complexity of $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}+N_{v}|d^{2}\right)\right)$, while the complexities of baseline methods based on RL and GNNs are $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}|d^{2}\right)\right)$. See Appendix D.6 for detailed explanations. While CONAL slightly increases the complexity compared to existing RL and GNN-based methods due to its heterogeneous modeling approach, it achieves significant performance improvements.
@@ 4 Experiments
In this section, we describe the experimental setup. To evaluate the effectiveness of CONAL, we compare it with several variations and state-of-the-art baselines in various network scenarios.
@@ 4.1 Experimental Settings
Simulation Configurations. Similar to most previous works (Geng et al., 2023; Yan et al., 2020), we evaluate the proposed framework in the simulation benchmarks that mimic various network systems. We adopt a Waxman topology with 100 nodes and nearly 500 links (Waxman, 1988) as the physical network, named WX100. Computing resources of physical nodes and bandwidth resources of physical links are uniformly distributed within the range of [50, 100] units. In default settings, for each simulation run, we create 1000 VN with varying sizes from 2 to 10. The computing resource demands of the nodes and the bandwidth requirements of the links within each VN are uniformly distributed within the range of [0, 20] and [0, 50] units, respectively. The virtual nodes in each VN are randomly interconnected with a probability of 50%. The lifetime of each VN follows an exponential distribution with an average of 500 time units. The arrival of these VNs follows a Poisson process with an average rate $\eta$, where $\eta$ denotes the average arrived VN count per unit of time. In the subsequent experiments, we manipulate the distribution settings of VNs and change the PN topologies to simulate various network systems.
@@
Compared Baselines. We compare CONAL with both heuristic and RL-based methods. The heuristic baselines includes node ranking-based methods (i.e., NRM-VNE (Zhang et al., 2018), GRC-VNE (Gong et al., 2014), NEA-VNE (Fan et al., 2023)) and meta-heuristics (i.e., GA-VNE (Zhang et al., 2019), PSO-VNE (Jiang & Zhang, 2021)). The learning-based baselines are PG-CNN (Ma et al., 2023), DDPG-ATT (He et al., 2023a), A3C-GCN (Zhang et al., 2023b), and GAL-VNE (Geng et al., 2023). See Appendix F.2 for their descriptions.
Evaluation Metrics. Following most previous research (Fischer et al., 2013; Yan et al., 2020), we evaluate the effectiveness of VNE algorithms with widely used key performance metrics: VN Acceptance Rate (VN_ACR); Long-Term REVenue (LT_REV); Long-Term Revenue-to-Consumption ratio (LT_R2C). We also consider AVeraGe Solving Time (AVG_ST) as an additional metric to measure the computational efficiency, due to the real-time demands of online network systems. Additionally, for CONAL and its variations, we employ Constraint VIOlation (C_VIO) to measure the degree of constraint satisfaction. See Appendix F.4 for their definitions.
@@ 4.2 Results and Analysis
Overall Evaluation. The results of VNE algorithms under default settings are shown in Table 1. We observe that CONAL outperforms baselines across three key performance metrics. This is due to CONAL’s ability to effectively perceive and handle the complex constraints of VNE. Among baselines, NEA-VNE, GA-VNE, and GAL-VNE are the best node-ranking-based heuristics, meta-heuristics, and RL-based methods. However, GA-VNE which relies on extensive search in the solution space has a longer running time, while GAL-VNE exhibits lower R2C metrics. Compared to NEA-VNE, GA-VNE and GAL-VNE, CONAL achieves improvements of (10.04%, 11.07%, 13.99%), (10.61%, 28.72%, 17.80%) and (4.77%, 24.04%, 6.21%) in term of (VN_ACR, LT_R2C, and LT_REV), respectively. We also observe that CONAL outperforms all RL-based baselines across key performance metrics, demonstrating the superiority of its modeling, optimization, and representation methods. Although CONAL’s running time is not the lowest, it remains competitive, comparable to NEA-VNE and A3C-GCN, and outperforms other baselines such as GA-VNE and MCTS-VNE. These results underscore the effectiveness of CONAL for VNE in providing high-quality and feasible solutions to improve resource utilization and request acceptance.
@@
Ablation Study. We design several variations to manifest the efficacy of each proposed component: $\text{CONAL}_{\text{w/o HM}}$, $\text{CONAL}_{\text{w/o PC}}$, $\text{CONAL}_{\text{w/o HM \& PC}}$, $\text{CONAL}_{\text{w/o REACH}}$, and $\text{CONAL}_{\text{w/o ARB}}$. See Appendix F.3 for their descriptions. The results of CONAL and its variations are presented in Table 1. Compared to the first three variants, CONAL achieves superior results across various performance metrics. This indicates that our constraint-aware graph representation method enhances the constraint awareness of policy. Notably, CONALw/o PC shows the most significant performance declines, even worse than CONALw/o HM & PC. This may be due to heterogeneous graph modeling methods increasing the link complexity of graph, which highlights the necessity for improving bandwidth constraint sensitivity. Furthermore, CONALw/o REACH and CONALw/o ARB also decrease performance, which demonstrates that our optimization method facilitates both resource utilization and constraint satisfaction. This study shows that each component of CONAL contributes to its overall performance, enhancing the its ability to handle and perceive the complex constraints of VNE.
Figure 3: Results in Training Stability Study.
Training Stability Study. To study the training stability of CONAL, we first compare the learning curves of CONAL and $\text{CONAL}_{\text{w/o ARB}}$, followed by an analysis of how different training conditions impact testing performance. (A) Learn curves analysis. We compare the learning curves of CONAL and $\text{CONAL}_{\text{w/o ARB}}$, depicted in Figure 3. We average the episode returns of 1000 VNs in one simulation as the metric, due to inherent differences of return resulting from VN size. We observe that while the performance of $\text{CONAL}_{\text{w/o ARB}}$ fluctuates during the training process, CONAL exhibits greater stability. This stability is attributed to our adaptive reachability budget method, which effectively addresses unsolvable instances during training, thus enhancing overall training stability. (B) Training Conditions vs. Testing Performance Analysis. Similar to our preliminary study, we further explore the impact of training conditions on CONAL performance. We train CONAL and A3C-GCN under different arrival rates $\lambda$ of VN requests from 0.14 to 0.26, as higher arrival rates increase the frequency of unsolvable instances. After training, we evaluate the models using a fixed arrival rate of $\lambda=0.14$ with a random seed of 0 as the benchmark. The results are shown in Figures 3(b)(c)(d). We observe that as the arrival rate $\lambda$ of training conditions increases, CONAL maintains a more stable testing performance while A3C-GCN shows an obvious decline. This analysis shows that CONAL is more stable in learning a high-quality policy by efficiently handling unsolvable instances and perceiving complex constraints.
@@
Generalizability Study. In practical network systems, fluctuations in traffic patterns and resource demands are inevitable due to varying service requirements and dynamic user behaviors. To study the generalizability, we test the trained CONAL model in various network conditions via the following experiments. See Appendix G.1 for details. (A) Request Frequency Sensitivity Study. We assess the sensitivity of CONAL to varying arrival rates of VN requests by adjusting the average request frequency $\eta$. As illustrated in Figure 5, CONAL consistently outperforms the baseline algorithms across all tested $\eta$ values, demonstrating its superior adaptability to changes in network traffic. This analysis highlights CONAL’s effectiveness in handling network scenarios with fluctuating request rates, ensuring stable performance even as competition for resources increases. (B) Dynamic Request Distribution Study. To simulate more realistic network conditions, we evaluate CONAL’s performance under varying VN request distributions by modifying resource demands and node sizes in different stages. As shown in Figure 6, CONAL demonstrates strong adaptability across all stages, outperforming the baseline algorithms even as the complexity of requests increases. This study shows CONAL’s ability to generalize effectively in dynamic networks with shifting requirements.
@@
Scalability Analysis. To assess the scalability of CONAL, we explore its performance in large-scale network systems and its time consumption to adapt to varying network sizes. See Appendix G.2 for details. In summary: (A) Large-scale Network Validation. We evaluate CONAL on a Waxman topology with 500 nodes and about 1300 links mimicking a large-scale cloud cluster. The results are shown in Figure 7. We observe that CONAL outperforms the baselines in most metrics in such a larger topology. This demonstrates the efficiency of CONAL in large-scale network scenarios. Results demonstrate that CONAL consistently outperforms all baseline models, even in this large-scale network system scenario. (B) Solving Time Scale Analysis. To investigate solving time scalability of CONAL, we increase the size of physical network from 200 to 1,000 nodes to simulate network systems of varying scales. The results, depicted in Figure 8, show that even at larger network scales, CONAL maintains efficient solving times while delivering excellent performance. This efficiency make CONAL a viable solution for real-time decision-making in large-scale network systems.
@@
Real-world network Validation. To verify the effectiveness of CONAL in real-world network systems, we conduct experiments on two well-known networks (Yan et al., 2020; He et al., 2023a): GEANT, a 40-node academic research network with 64 edges, and BRAIN, a 161-node high-speed data network with 166 edges. The results shown in Table 3 reveals that CONAL outperforms all baselines across both network systems in terms of performance metrics. This validation shows the effectiveness of CONAL in practical systems and various topologies. See Appendix G.3 for details.
Hyperparameter Sensitivity Study. We explore the impact of the following two key hyperparameters on performance. (A) Update interval $\mu$ of surrogate policy. The results are shown in Figure 9. We observe that extremely frequent updates potentially lead to instability and divergence in the learning process. Additionally, too slow updates do not offer significant further benefits and may even increase computational overhead. (B) Augment ratio $\epsilon$ used in the path-bandwidth contrast module. The results shown in Figure 10 reveal that a reasonable augment ratio $\epsilon$ enhances the model’s sensitivity to bandwidth constraints. However, excessively high values of $\epsilon$ yield minimal improvements or may even harm performance. See Appendix G.4 for details.
@@ 5 Conclusion
In this paper, we proposed the CONAL for VNE to enhance constraint management and training robustness, which is critical to the performance and reliability of network systems. Specifically, we formulated the VNE problem as a violation-tolerant CMDP to optimize both the quality and feasibility of solutions. This method allows us to always obtain complete solutions to precisely evaluate the quality of the solution. Additionally, we presented a reachability-guided optimization objective with an adaptive feasibility budget method to ensure persistent constraint satisfaction while alleviating the conservativeness of policy. This approach also address the instability of policy optimization caused by unsolvable instances. Furthermore, to finely perceive the complex constraints of VNE, we proposed a constraint-aware graph representation method, which consists of a heterogeneous modeling module for indicating cross-graph relations and a path-bandwidth contrast module that enhances the sensitivity to bandwidth constraints. Finally, we conducted extensive experiments to verify the effectiveness of our proposed methods. In the future, we plan to tackle additional constraints in specific networking scenarios, e.g., latency-aware edge computing and energy-efficient green computing. This will require specialized design efforts to address scenario-specific challenges effectively.
@@ Appendix A Related Work
VNE Algorithms. To solve this challenging and significant problem, many approaches have been designed for VNE, which can be classified as exact, heuristics, and learning-based methods. Initially, exact algorithms formulate the VNE problem as integer linear programming (Shahriar et al., 2018) or mixed integer linear programming (Chowdhury et al., 2009), and then solve them with exact solvers. However, this is impractical due to extensive computations and time consumption. Thus, numerous heuristic algorithms have been proposed to offer solutions within an acceptable time, such as node ranking strategies (Zhang et al., 2018; Gong et al., 2014; Fan et al., 2023), meta-heuristics (Dehury & Sahoo, 2019; Zhang et al., 2019; Jiang & Zhang, 2021), etc. However, they heavily rely on manual heuristic design and merely for specific scenarios. Recently, reinforcement learning has emerged as a promising solution for VNE and many RL-based VNE algorithms have been proposed (Haeri & Trajković, 2017; Wang et al., 2021; Zhang et al., 2022; He et al., 2023a; Zhang et al., 2023b; Geng et al., 2023). In general, they model the solution construction process of each VNE instance as an MDP, but do not consider the fine-grained constraint violations. Then, they leverage existing network networks (e.g., convolutional neural network, GNN, etc.) to extract features from PN and VN, separately. Finally, they optimize the model with different RL methods (e.g., asynchronous advantage actor-critic, PPO, etc.). However, they struggle to handle such complex constraints of VNE thereby compromising performance, since they lack constraint awareness in MDP modeling, representation learning, and policy optimization.
@@
RL for COPs. The application of RL to solve COPs has emerged as a hot topic in decision-making, which aims to learn efficient solving strategies from data (Bengio et al., 2021). Many efforts have been directed toward classic COPs such as routing (Zhou et al., 2023; Ye et al., 2023), scheduling (Zhang et al., 2023a; 2020), bin packing (Pan et al., 2023; Zhao et al., 2021), etc. These approaches can be broadly categorized into two types based on their solving processes: construction and improvement. While improvement methods start with an initial solution and use an RL policy to iteratively refine it, construction methods build a solution incrementally from scratch (Mazyavkina et al., 2021). Construction methods typically employ RL to guide the sequential selection to form a complete solution. Given the real-time requirements of practical network systems, most existing RL-based VNE algorithms are designed as construction methods to provide solutions within an acceptable time (Zhang et al., 2022; He et al., 2023a; Zhang et al., 2023b). In contrast to these classic COPs, VNE presents unique complexities in representation learning due to its multifaceted and hard constraints, such as the interaction of cross-graph status and the assessment of bandwidth-constrained path connectivity. Furthermore, the existence of unsolvable instances in the training process can compromise robustness, potentially causing ineffective policies.
@@
Safe RL Methods. Safe RL aims to maximize the expected rewards while ensuring safety constraints are not violated (Gu et al., 2022). Early efforts in safe RL focus on keeping cumulative constraint violations below a fixed cost budget (Achiam et al., 2017; Ray et al., 2019; Yang et al., 2022). To address the stricter constraint requirements of practical applications, recent works have proposed achieving state-wise safety, which ensures the satisfaction of instantaneous constraints at each decision timestep (Zhao et al., 2023; 2022; He et al., 2023b; Yu et al., 2022). One promising approach involves incorporating reachability analysis into the CMDP framework (Yu et al., 2022). This method employs a reachability function to assess the state feasibility, which significantly expands the feasible set of policy, and mitigates the conservativeness of the policy. However, modeling VNE as a reachability-guided CMDP presents challenges. Concretely, In the process of solution construction, if any constraints are violated, the process will be early terminated and lead to incomplete solutions. This hinders precise measurement of both the quality of solutions and the degree of constraint violations. Additionally, these methods typically assume a non-empty feasible set. But for VNE, where it is hard to avoid facing unsolvable instances without any feasible solution, these approaches often confront the challenges of unstable policy optimization, since the constraints of these instances are impossibly satisfied.

### v2

@@ 3.3 Constraint-aware Graph Representation
The VNE processing is governed by complex and multifaceted constraints, presenting challenges in representation learning. These include the interaction of cross-graph status and the assessment of bandwidth-constrained path connectivity. To address this issue, we propose a constraint-aware graph representation method with a heterogeneous modeling module for cross-graph status fusion, and a contrastive learning-based module to enhance path connectivity awareness. It efficiently perceives the complex constraints of VNE, providing a higher guarantee of solution feasibility.
@@
To address this challenge, we propose a novel path-bandwidth contrast method to enhance bandwidth constraint awareness via contrastive learning, whose core idea is creating augmented views with feasibility-consistency augmentations and making node representations in these views close.
Definition 3.5 (Feasibility-consistency Augmentations).
@@
These augmentations generate multiple views of the original heterogeneous graph, which maintain the same feasibility semantics before and after their application. Following this principle, we develop several augmentation methods by modifying the topology of either VN or PN without impacting solution feasibility. (a) Physical Link Addition $\phi^{A}$. We add a specific number $\epsilon\cdot|N^{p}|$ of physical links in PN, whose bandwidth resources are equal to the difference between the smallest requirements among all virtual links and 1, i.e., $\min_{l_{v}\in L_{v}}\mathcal{B}(l_{v})-1$. (b) Virtual Link Addition $\phi^{B}$. We add a specific number $\epsilon\cdot|N^{v}|$ of virtual links that require a zero bandwidth resource to enhance the complexity and connectivity of the VN. Here, $\epsilon$ is an augment ratio that determines the proportion of links to be added based on the number of nodes in the network.
After applying these augmentations, we create two new views $G_{I}^{A}=\phi_{A}(G_{I})$ and $G_{I}^{B}=\phi^{B}(G_{I})$, which have same feasibility semantics of VNE instance $I$. Using our heterogeneous graph network, we extract node representations under views $G_{I}^{A}$ and $G_{I}^{B}$, denoted as $Z^{A}$ and $Z^{B}$, respectively. Subsequently, we utilize contrastive learning to enhance the proximity of node representations under the augmented views $G_{I}^{A}$ and $G_{I}^{B}$. This necessitates that the model precisely discerns the noisy implications of those links with less bandwidth that play no impact on solution feasibility. Through this method, we aim to enhance the model’s sensitivity towards link bandwidth, effectively mitigate the influence of irrelevant links in the GNN propagation process, and bolster its overall awareness of bandwidth constraints. In this work, we adopt the Barlow Twins method (Zbontar et al., 2021) for its simplicity and effectiveness, which circumvents negative sample selection and maintains the original network architecture. Given the embeddings under two augmented views, $H^{a}$ and $H^{b}$, we use this contrastive loss $L_{\textit{\tiny CL}}$ to reduce redundancy between embedding components by aligning their cross-correlation matrix with the identity matrix. This unsupervised loss can be seamlessly integrated into the training process of RL. See Appendix D.4 for the details of Barlow Twins method.
@@ 3.4 Computational Complexity Analysis
Note that CONAL solely uses its surrogate policy and path-bandwidth contrast module for policy optimization during training. During inference, CONAL has a computational complexity of $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}+N_{v}|d^{2}\right)\right)$. See Appendix D.6 for detailed explanations.
@@ 4 Experiments
To evaluate the effectiveness of CONAL, we compare it with state-of-the-art baselines in various network scenarios.
@@ 4.1 Experimental Settings
Simulation Configurations. Similar to most previous works (Geng et al., 2023; Yan et al., 2020), we evaluate CONAL in Virne benchmarks that simulate various network systems (See Appenidx H for its introduction). We adopt a widely-used Waxman topology with 100 nodes and nearly 500 links (Waxman, 1988) as the physical network, named WX100. Computing resources of physical nodes and bandwidth resources of physical links are uniformly distributed within the range of [50, 100] units. In default settings, for each simulation run, we create 1000 VN with varying sizes from 2 to 10. The computing resource demands of the nodes and the bandwidth requirements of the links within each VN are uniformly distributed within the range of [0, 20] and [0, 50] units, respectively. The virtual nodes in each VN are randomly interconnected with a probability of 50%. The lifetime of each VN follows an exponential distribution with an average of 500 time units. The arrival of these VNs follows a Poisson process with an average rate $\eta$, where $\eta$ denotes the average arrived VN count per unit of time. In the subsequent experiments, we manipulate the distribution settings of VNs and change the PN topologies to simulate various network systems.
@@
Compared Baselines. We compare CONAL with both heuristic and RL-based methods. The heuristic baselines includes node ranking-based methods (i.e., NRM-VNE (Zhang et al., 2018), GRC-VNE (Gong et al., 2014), NEA-VNE (Fan et al., 2023)) and meta-heuristics (i.e., GA-VNE (Zhang et al., 2019), PSO-VNE (Jiang & Zhang, 2021)). The learning-based baselines are PG-CNN (Ma et al., 2023), DDPG-ATT (He et al., 2023a), A3C-GCN (Zhang et al., 2024b), and GAL-VNE (Geng et al., 2023). See Appendix F.2 for their descriptions.
Evaluation Metrics. Following most previous research (Fischer et al., 2013; Yan et al., 2020), we adopt widely used key metrics for VNE algorithm evaluation: VN Acceptance Rate (VN_ACR); Long-Term REVenue (LT_REV); Long-Term Revenue-to-Consumption ratio (LT_R2C). We also consider AVeraGe Solving Time (AVG_ST) as an additional metric to measure the computational efficiency, due to the real-time demands of online network systems. Additionally, for CONAL and its variations, we employ Constraint VIOlation (C_VIO) to measure the degree of constraint satisfaction. See Appendix F.4 for their definitions.
@@ 4.2 Results and Analysis
Overall Evaluation. The results of VNE algorithms under default settings are shown in Table 1. We observe that CONAL outperforms baselines across three key performance metrics. This is due to CONAL’s ability to effectively perceive and handle the complex constraints of VNE. Among baselines, NEA-VNE, GA-VNE, and GAL-VNE are the best node-ranking-based heuristics, meta-heuristics, and RL-based methods. However, GA-VNE which relies on extensive search in the solution space has a longer running time, while GAL-VNE exhibits lower R2C metrics. Compared to NEA-VNE, GA-VNE and GAL-VNE, CONAL achieves improvements of (10.04%, 11.07%, 13.99%), (10.61%, 28.72%, 17.80%) and (4.77%, 24.04%, 6.21%) in term of (VN_ACR, LT_R2C, and LT_REV), respectively. We also observe that CONAL outperforms all RL-based baselines across key performance metrics, demonstrating the superiority of its modeling, optimization, and representation methods. Although CONAL’s running time is not the lowest, it remains competitive, comparable to NEA-VNE and A3C-GCN, and outperforms other baselines such as GA-VNE and MCTS-VNE. These results underscore that CONAL provides high-quality and feasible solutions to improve resource utilization and request acceptance.
@@
Ablation Study. We design several variations to manifest the efficacy of each proposed component: $\text{CONAL}_{\text{w/o HM}}$, ${}_{\text{w/o PC}}$, ${}_{\text{w/o HM \& PC}}$, ${}_{\text{w/o REACH}}$, and ${}_{\text{w/o ARB}}$. See Appendix F.3 for their descriptions. The results are presented in Table 1. Compared to the first three variants, CONAL achieves superior results across various performance metrics. This indicates that our constraint-aware graph representation method enhances the constraint awareness of policy. Notably, CONALw/o PC shows the most significant performance declines, even worse than CONALw/o HM & PC. This may be due to heterogeneous graph modeling methods increasing the link complexity of graph, which highlights the necessity for improving bandwidth constraint sensitivity. Furthermore, CONALw/o REACH and CONALw/o ARB also decrease performance, which demonstrates that our optimization method facilitates both resource utilization and constraint satisfaction. This study shows that each component of CONAL contributes to its overall performance, enhancing its ability to handle and perceive the complex constraints of VNE.
[…]
@@
Generalizability Study. Practical network systems face fluctuations in traffic patterns and resource demands due to dynamic user service requirements. To study generalizability, we test trained models in various network conditions. (A) Request Frequency Sensitivity Study. We assess the sensitivity of CONAL to varying arrival rates of VN requests by adjusting the average request frequency $\eta$. Results illustrated in Figure 6 highlight CONAL’s effectiveness in handling network scenarios with fluctuating request rates, even as competition for resources increases. (B) Dynamic Request Distribution Study. We simulate varying network conditions by modifying resource demands and node sizes of VN requests in different stages. Results in Figure 7 show that CONAL can generalize effectively in dynamic systems with shifting requirements. See Appendix G.2 for details.
@@
Scalability Analysis. To assess the scalability of CONAL, we explore its performance in large-scale network systems and its time consumption to adapt to varying network sizes. In summary: (A) Large-scale Network Validation. We evaluate CONAL on a Waxman topology with 500 nodes representing a large-scale cloud cluster. The results demonstrate that CONAL consistently outperforms all baseline models, even in large-scale network systems. (B) Solving Time Scale Analysis. We increase the size of physical network from 200 to 1,000 nodes to simulate network systems of varying scales. The results show that even at larger network scales, CONAL maintains efficient solving times while delivering excellent performance. See Appendix G.3 for details.
@@
Real-world Network Topology Validation. To verify the effectiveness of CONAL in real-world network systems, we conduct experiments on two well-known networks (Yan et al., 2020; He et al., 2023a): GEANT and BRAIN. The results shown in Table 3 reveal that CONAL outperforms all baselines across both network systems on performance, which shows the effectiveness of CONAL in practical systems and various topologies. See Appendix G.4 for details.
Extension to Latency-aware VNE. We further study the VNE variation in edge computing with latency consideration, which additionally introduces such constraint. Results in Table 4 show that CONAL exhibits superior performance which efficiently handles complex multi-type constraints in this variation of VNE. See Appendix G.5 for details.
@@ 5 Conclusion
In this paper, we proposed the CONAL for VNE to enhance constraint management and training stability, which is critical to the performance and reliability of network systems. Specifically, we formulated the VNE problem as a violation-tolerant CMDP to optimize both the quality and feasibility of solutions, which precisely evaluates the quality and violation of the solution. Additionally, we presented a reachability-guided optimization objective with an adaptive feasibility budget method to ensure persistent constraint satisfaction while alleviating the conservativeness of policy. This approach also addresses the instability of policy optimization caused by unsolvable instances. Furthermore, to finely perceive the complex constraints of VNE, we proposed a constraint-aware graph representation method, consisting of a heterogeneous modeling module and a path-bandwidth contrast module. Finally, we conducted extensive experiments to verify the effectiveness of CONAL.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning for Combinatorial Optimization and Networking Resource Allocation. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
[…]
@@ Appendix A Related Work
VNE Algorithms. To solve this challenging and significant problem, many approaches have been designed for VNE, which can be classified as exact, heuristics, and learning-based methods. Initially, exact algorithms formulate the VNE problem as integer linear programming (Shahriar et al., 2018) or mixed integer linear programming (Chowdhury et al., 2009), and then solve them with exact solvers. However, this is impractical due to extensive computations and time consumption. Thus, numerous heuristic algorithms have been proposed to offer solutions within an acceptable time, such as node ranking strategies (Zhang et al., 2018; Gong et al., 2014; Fan et al., 2023), meta-heuristics (Dehury & Sahoo, 2019; Zhang et al., 2019; Jiang & Zhang, 2021), etc. However, they heavily rely on manual heuristic design and merely for specific scenarios. Recently, reinforcement learning has emerged as a promising solution for VNE and many RL-based VNE algorithms have been proposed (Haeri & Trajković, 2017; Wang et al., 2021; Zhang et al., 2022; He et al., 2023a; Zhang et al., 2024b; Geng et al., 2023; Wang et al., 2024). In general, they model the solution construction process of each VNE instance as an MDP, but do not consider the fine-grained constraint violations. Then, they leverage existing network networks (e.g., convolutional neural network, GNN, etc.) to extract features from PN and VN, separately. Finally, they optimize the model with different RL methods (e.g., asynchronous advantage actor-critic, PPO, etc.). Although they use the action masking mechanism to select physical nodes with adequate computing resources and avoid unnecessary failures, they can not address potential violations of path-level bandwidth constraints during state transitions, which still lead to numerous failure samples during exploration. Consequently, they struggle to handle such complex constraints of VNE thereby compromising performance, since they lack constraint awareness in MDP modeling, representation learning, and policy optimization.
@@
RL for COPs. The application of RL to solve COPs has emerged as a hot topic in decision-making, which aims to learn efficient solving strategies from data (Bengio et al., 2021). Many efforts have been directed toward classic COPs such as routing (Ye et al., 2023; Bi et al., 2024), scheduling (Zhang et al., 2023, 2024a), bin packing (Pan et al., 2023; Zhao et al., 2021), etc. These approaches can be broadly categorized into two types based on their solving processes: construction and improvement. While improvement methods start with an initial solution and use an RL policy to iteratively refine it, construction methods build a solution incrementally from scratch (Mazyavkina et al., 2021). Construction methods typically employ RL to guide the sequential selection to form a complete solution. Given the real-time requirements of practical network systems, most existing RL-based VNE algorithms are designed as construction methods to provide solutions within an acceptable time (Zhang et al., 2022; He et al., 2023a; Zhang et al., 2024b). In contrast to these classic COPs, VNE presents unique complexities in representation learning due to its multifaceted and hard constraints, such as the interaction of cross-graph status and the assessment of bandwidth-constrained path connectivity. Furthermore, the existence of unsolvable instances in the training process can compromise robustness, potentially causing ineffective policies.
@@
Safe RL Methods. Safe RL aims to maximize the expected rewards while ensuring safety constraints are not violated (Gu et al., 2022). Early efforts in safe RL focus on keeping cumulative constraint violations below a fixed cost budget (Achiam et al., 2017; Ray et al., 2019; Yang et al., 2022). To address the stricter constraint requirements of practical applications, recent works have proposed achieving state-wise safety, which ensures the satisfaction of instantaneous constraints at each decision timestep (Zhao et al., 2023, 2022; He et al., 2023b; Yu et al., 2022). One promising approach involves incorporating reachability analysis into the CMDP framework (Yu et al., 2022). This method employs a reachability function to assess the state feasibility, which significantly expands the feasible set of policy, and mitigates the conservativeness of the policy. However, modeling VNE as a reachability-guided CMDP presents challenges. Concretely, In the process of solution construction, if any constraints are violated, the process will be early terminated and lead to incomplete solutions. This hinders precise measurement of both the quality of solutions and the degree of constraint violations. Additionally, these methods typically assume a non-empty feasible set. But for VNE, where it is hard to avoid facing unsolvable instances without any feasible solution, these approaches often confront the challenges of unstable policy optimization, since the constraints of these instances are impossibly satisfied.